# Seamless Monitoring of Stress Levels Leveraging a Universal Model for Time Sequences

### v1

@@ Abstract
Monitoring the stress level in patients with neurodegenerative diseases can help manage symptoms, improve patient’s quality of life, and provide insight into disease progression. In the literature, ECG, actigraphy, speech, voice, and facial analysis have proven effective at detecting patients’ emotions. On the other hand, these tools are invasive and do not integrate smoothly into the patient’s daily life. HRV has also been proven to effectively indicate stress conditions, especially in combination with other signals. However, when HRV is derived from less invasive devices than the ECG, like smartwatches and bracelets, the quality of measurements significantly degrades. This paper presents a methodology for stress detection from a smartwatch based on a universal model for time series, UniTS, which we fine-tuned for the task. We cast the problem as anomaly detection rather than classification to favor model adaptation to individual patients and allow the clinician to maintain greater control over the system’s predictions. We demonstrate that our proposed model considerably surpasses 12 top-performing methods on 3 benchmark datasets. Furthermore, unlike other state-of-the-art systems, UniTS enables seamless monitoring, as it shows comparable performance when using signals from invasive or lightweight devices.
Keywords Anomaly Detection  $\cdot$ Seamless Monitoring  $\cdot$ Foundation Models  $\cdot$ Transformers  $\cdot$ Neurodegenerative Diseases  $\cdot$ Stress Detection
@@ 1 Introduction
High stress levels can worsen the symptoms of neurodegenerative diseases such as Alzheimer’s and Parkinson’s [22]. In these patients, stress is generated by mood changes such as depression and apathy, confinement at home, and other factors such as social isolation, uncertainty about the evolution of the disease, and the economic, health, personal, and family situation [23] [18]. Analyzing stress levels can provide insights into disease progression and help healthcare providers adjust treatment plans accordingly. Managing stress can also improve patients’ quality of life and help to inform new therapeutic strategies.
AI-based methods for stress detection are motivated by the need for early, timely, and personalized stress management solutions. These methods leverage cutting-edge machine learning algorithms to provide objective, real-time monitoring and analysis, significantly contributing to the field of mental health and well-being (see, among others, [40]).
Among the numerous techniques presented in the literature, those based on data extracted from sensor devices are particularly promising due to several advantages. One of the key benefits of using sensor-based methods is their ability to provide real-time data, which is crucial for timely interventions and stress management. Secondly, they protect privacy more than solutions based on behavioral signals such as speech, gestures, and facial expressions [26], a particularly relevant aspect for elderly and frail people. Third, they may support continuous monitoring, enabling more convenient home care solutions for patients and caregivers.
In the literature, various sensor types are used to detect stress, leveraging different physiological signals. As surveyed in [36], commonly used devices are ECG and EDA, EEG, skin temperature, respiratory and pressure sensors, and activity sensors like accelerometers, gyroscopes, and RFID technology. Among these devices, only a few are well-suited for seamless monitoring due to factors such as invasiveness, lack of autonomy for patients, and high costs. We define seamless monitoring as the real-time collection of signals that is i) without time limits, ii) uninterrupted, and iii) does not require explicit actions by patients and/or doctors [^1] In this sense, ECG is not seamless, because it can usually last for 24-48 hours, and must be activated by the doctor or patient, depending upon the device type.. In other terms, while continuous monitoring only ensures that data is captured without interruption, seamless monitoring ensures that this process integrates smoothly into the patient’s daily life. To address this challenge, which is essential to facilitate the widespread adoption of monitoring technologies for stress detection and healthcare applications in general, we propose a methodology providing the following contributions:
@@
    To detect stress signals from sensor data, we leverage for the first time a universal model for time sequences, UniTS [12], achieving 9% superior performance against the best performing – i.e., MSCRED [39] – among 12 compared state-of-the-art (SoTA) methods on 3 publicly available datasets;
@@
    We cast the problem of stress detection as anomaly detection rather than classification, as the majority of works in the literature do (see Table 1). Anomaly detection typically involves establishing a baseline of normality rather than learning boundaries between multiple classes. The former is a more straightforward process since any deviation from normal behavior can be visualized and investigated for specific reasons. This way, doctors can more easily understand and trust the system’s predictions, thereby enhancing their ability to provide high-quality care.
@@
    We release our code publicly [^2] https://github.com/davegabe/Wearable-Stress-Monitor with all the model weights to encourage reproducibility from future researchers.
@@ 2 Related Work
Our work relates to the fields of AI-based emotion monitoring from sensor data and anomaly detection from time series.
@@
[44] ECG S ECG Wearable Conv. Gaussian Mixture VAE AD 2023
@@ 2.1 Emotion Monitoring and Stress Detection
Table 1 summarizes the key details regarding research on emotion recognition through physiological signals, providing a qualitative comparison with the solution proposed in this paper. (1) “Signals Used” lists the physiological signals employed. Examples are Electrocardiogram (ECG), Heart Rate (HR), Galvanic Skin Response (GSR), Respiration (Resp.), and Electromyogram (EMG), among others. The choice of signals is crucial as it directly affects the accuracy and reliability of the detection system. Different signals provide varied information about an individual’s physiological state, impacting the performance of the employed models. (2) “Application” specifies whether the study focuses on stress (S) or emotion (E) detection, whereas the second considers more nuanced mental conditions. (3) “Device Type” specifies the type of device used to collect the physiological signals, such as wearables, smartwatches, wristbands, GSR devices, or pulse oximetry. The device type influences the monitoring system’s accuracy, practicality, comfort, and user acceptance. Only some types of devices allow for seamless monitoring in real-world settings, which is especially valuable for frail people. (4) “Model Type” lists the machine learning or statistical models used for classification tasks. Notice how, with the exception of [2] and [44], all works rely on simple off-the-shelf machine learning models (SVM, kNN, and pLDA [^3] Extended Linear Discriminant Analysis.), which might not be suitable to detect emotions or stress levels that are not in the training data distribution, and therefore have limited capacity for generalization but also for personalization. In [2], an ensemble mechanism over Gradient Boosts (GB) is applied to classify stress levels by leveraging the expertise of multiple models. In this case, a linear regression meta-model is applied to the output of the GB base models to predict the final outcome. In [44], a clustering algorithm based on a Gaussian mixture convolutional variational autoencoder is employed to identify normal emotional patterns and detect anomalies. To the best of our knowledge, we are the first to use a state-of-the-art Transformer model to tackle stress and abnormal emotion detection. (5) “Task” indicates whether the model’s task is classification (C) or anomaly detection (AD). Most works tackle stress and emotion monitoring as classification problems, categorizing states into predefined classes. The only exception is UAED [44], which performs an anomaly detection task. We argue that approaching stress monitoring as an anomaly detection problem is more principled and inherently more explainable than doing fine-grained emotion classification.
@@ 2.2 Anomaly Detection in Time Series
Following [44], we cast stress monitoring as a problem of anomaly detection. We identified five categories of methods proposed in the literature for anomaly detection in time series. Distance-based outlier detectors consider the distance of a point from its k-nearest neighbors. Density-based methods (among others [38, 35]) consider the density of the point and its neighbors. Prediction-based methods (among others, [5]) calculate the difference between the predicted and true values to detect anomalies. Reconstruction-based methods (among others [11]) compare the input signal and the reconstructed one in the output layer, typically using autoencoders. These methods assume anomalies are difficult to reconstruct and lost when the signal is mapped to lower dimensions, so a higher reconstruction error indicates a higher anomaly score.
Recently, GANs have been employed for anomaly detection in time series data. MAD-GAN [25] combines the discriminator output and reconstruction error to detect anomalies in multivariate time series. BeatGAN [43] uses an encoder-decoder generator with a modified time-warping-based data augmentation to detect anomalies in medical ECG inputs. TadGAN [13] uses a cycle-consistent GAN architecture with an encoder-decoder generator and proposes several ways to compute reconstruction error combined with critic outputs. HypAD [11] proposes using hyperbolic uncertainty for anomaly detection. Tuli et al. [37] introduce TranAD, built on deep transformer networks. This model leverages attention-based sequence encoders to infer broader temporal patterns within the data. TranAD employs focus score-based self-conditioning to enhance its performance for effective multi-modal feature extraction and utilizes adversarial training to ensure stability. Finally, Zhu et al. [44] rely on a 2d Convolutional Variational Autoencoder with a Gaussian Mixture Model as the latent representation distribution. Interestingly, the authors extrapolate ECG signals, stack them vertically on top of one another, and process them as images to detect anomalies unsupervisedly.
@@ 3 Method
Figure 1: Finetuning of the UniTS architecture for anomaly detection. The multivariate time series in input gets transformed into several tokens, which are then passed through $N$ different blocks of the UniTS model. UniTS reconstructs the sample tokens, which are unpatched and compared against the real tokens. The comparison uses a dynamic threshold to discern between anomalies and normality. For visualization purposes, we do not illustrate the prompt tokens concatenated to the sample ones.
UniTS [12] is the first foundation model444foundation models are pre-trained on broad data at scale and can be adapted to a wide range of downstream tasks. for multivariate time series, designed to perform in a unified way various tasks such as forecasting, imputation, classification, and anomaly detection. It employs self-attention to capture relationships across both the sequence and variable dimensions. This is particularly beneficial in our multi-sensor domain since if one variable has an anomaly, the model can assess how this anomaly correlates with the patterns observed in other variables, thereby improving detection accuracy. Furthermore, UniTS is designed to integrate signals of different types and lengths in a cohesive manner. To achieve this, it employs multi-scale processing techniques to handle time-series data of different lengths, which involves processing the data at various temporal resolutions.
@@ 3.1 Problem Definition and Preliminaries
For each time point $t$, an anomaly score $S_{t}=g(\mathbf{x}_{t},f(\mathbf{X}_{train}))$ is computed, where $g$ measures the deviation of $\mathbf{x}_{t}$ from the expected behavior defined by $f$. An observation $\mathbf{x}_{t}$ is classified as anomalous if its anomaly score $S_{t}$ exceeds a predefined threshold $\tau$. Thus, the set of detected anomalies is given by $\mathbf{A}=\{\mathbf{x}_{t}\mid S_{t}>\tau,t\in\{1,2,\ldots,T\}\}$.
@@ 3.2 Framework Overview
UniTS – see Figure 1 – is a versatile multi-task model featuring a unified network architecture. We exploit the original architecture to perform anomaly detection in time series data, as described hereafter.
Input Tokenization. UniTS employs a token-based approach to represent tasks and time series from various domains. It introduces two [^5] The original paper proposes three different tokens, among which also the task tokens, which we do not use in our scenario. We point the reader to [12] for more details on how task tokens are used. unique token types: i.e., sample, and prompt, each fulfilling a specific role in time series analysis.
@@
- Prompt tokens. They are defined as learnable embeddings $z_{p}\in\mathbb{R}^{p\times n\times d}$ where $p$ is the number of tokens. These tokens incorporate the task UniTS needs to perform. Notice that, in Figure 1, we do not illustrate the prompt tokens proposed in the original paper since we only exploit the forecasting prompt in our anomaly detection scenario. Nevertheless, for the sake of completeness, we invite the reader to visualize that these prompt tokens are prepended to the input series tokens as in Eq. (2).
@@
UniTS modules. The model takes in input $z_{\text{Anomaly}}$ and feeds it to $N$ blocks of modified transformer architecture to handle heterogeneous multi-domain data with varying dynamics and the number of variables. We use the original UniTS architecture. In more detail, we use Time and Variable Multi-Head Self-Attention (MHSA) blocks, a Dynamic Feed-Forward Network (FFN), and gating modules. We refer the reader to the original paper’s appendix [12] for more detail on the transformer and tower modules.
- Time and Variable MHSA. As suggested in the original paper, we use a two-way self-attention for feature and time dimensions. This approach contrasts with previous methods that apply self-attention to either time or variable dimension but not to both dimensions. Time and variable self-attention effectively handle time series samples with various numbers of features $n$ and different time lengths $t$.
- Dynamic FFN. The transformer block is modified by incorporating a dynamic operator into an FFN layer. This modification enables the FFN to capture dependencies between tokens, in contrast to the standard FFN, which processes embedding vectors on a point-wise basis.

### v2

# Seamless Monitoring of Stress Levels Leveraging a Foundational Model for Time Sequences

@@ Abstract
Monitoring the stress level in patients with neurodegenerative diseases can help manage symptoms, improve patient’s quality of life, and provide insight into disease progression. In the literature, ECG, actigraphy, speech, voice, and facial analysis have proven effective at detecting patients’ emotions. On the other hand, these tools are invasive and do not integrate smoothly into the patient’s daily life. HRV has also been proven to effectively indicate stress conditions, especially in combination with other signals. However, when HRV is derived from less invasive devices than the ECG, like wristbands and smartwatches, the quality of measurements significantly degrades. This paper presents a methodology for stress detection from a wristband based on a universal model for time series, UniTS, which we finetuned for the task and equipped with explainability features. We cast the problem as anomaly detection rather than classification to favor model adaptation to individual patients and allow the clinician to maintain greater control over the system’s predictions. We demonstrate that our proposed model considerably surpasses 12 top-performing methods on three benchmark datasets. Furthermore, unlike other state-of-the-art systems, UniTS enables seamless monitoring, as it shows comparable performance when using signals from invasive or lightweight devices.
keywords:
Anomaly Detection , Seamless Monitoring , Foundation Models , Transformers , Neurodegenerative Diseases , Stress Detection , Explainability
journal: Journal of Artificial Intelligence in Medicine
[…]
[sapienza] organization=Sapienza University of Rome, addressline=Via Salaria 113, city=Rome, postcode=00198, country=Italy
[…]
[tum] organization=Technical University of Munich, addressline=Richard-Wagner-Str. 1, city=Munich, postcode=80333, country=Germany
@@ 1 Introduction
High stress levels can worsen the symptoms of neurodegenerative diseases such as Alzheimer’s and Parkinson’s [25]. In these patients, stress is generated by mood changes such as depression and apathy, confinement at home, and other factors such as social isolation, uncertainty about the evolution of the disease, and the economic, health, personal, and family situation [26] [21]. Analyzing stress levels can provide insight into disease progression and help healthcare providers adjust treatment plans accordingly. Managing stress can also improve patients’ quality of life and help to inform new therapeutic strategies.
AI-based methods for stress detection are motivated by the need for early, timely, and personalized stress management solutions. These methods leverage cutting-edge machine learning algorithms to provide objective, real-time monitoring and analysis, significantly contributing to mental health and well-being (see, among others, [45]).
Among the numerous techniques presented in the literature, those based on data extracted from sensor devices are particularly promising due to several advantages. One of the key benefits of using sensor-based methods is their ability to provide real-time data, which is crucial for timely interventions and stress management. Secondly, they protect privacy more than solutions based on behavioral signals such as speech, gestures, and facial expressions [29], a particularly relevant aspect for elderly and frail people (see [32], among others). Third, they may support continuous monitoring, enabling more convenient home care solutions for patients and caregivers.
In the literature, various sensor types are used to detect stress, leveraging different physiological signals. As surveyed in  [39], commonly used devices are ECG and EDA, EEG, skin temperature, respiratory and pressure sensors, and activity sensors like accelerometers, gyroscopes, and RFID technology. Among these devices, only a few are well-suited for seamless monitoring due to invasiveness, lack of autonomy for patients, and high costs. We define seamless monitoring as the real-time collection of signals that is i) without time limits, ii) uninterrupted, and iii) does not require explicit actions by patients and doctors. [^3] In this sense, ECG is not seamless because it can usually last for 24-48 hours, and must be activated by the doctor or patient, depending upon the device type. In other terms, while continuous monitoring only ensures that data is captured without interruption, seamless monitoring ensures that this process integrates smoothly into the patient’s daily life. To address this challenge, which is essential to facilitate the widespread adoption of monitoring technologies for stress detection and healthcare applications in general, we propose a methodology providing the following contributions:
@@
    To detect stress signals from sensor data, we leverage for the first time a universal model for time sequences, UniTS [14], achieving 22% superior performance against the second best performing – i.e., DAGMM [49] – among 12 compared state-of-the-art (SoTA) methods on three publicly available datasets;
@@
    We cast the problem of stress detection as anomaly detection rather than classification, as the majority of works in the literature do (see Table 1). Anomaly detection typically involves establishing a baseline of normality rather than learning boundaries between multiple classes. The former is a more straightforward process since any deviation from normal behavior can be visualized and investigated for specific reasons. This way, doctors can more easily understand and trust the system’s predictions, enhancing their ability to provide high-quality care.
@@
    We integrate an LLM to translate anomalies in human-readable explanations based on personalized prompting [43] to support healthcare personnel in interpreting the detections UniTS makes.
  5. 5.
    We release our code publicly with all the model weights to encourage reproducibility from future researchers. [^4] https://github.com/davegabe/Wearable-Stress-Monitor
@@ 2 Related Work
Our work relates to AI-based emotion monitoring from sensor data and anomaly detection from time series.
@@
[48] ECG S ECG, Wearable Conv. Gaussian Mixture VAE AD 2023
@@ 2.1 Emotion Monitoring and Stress Detection
Table 1 summarizes the key details regarding research on emotion recognition through physiological signals, providing a qualitative comparison with the solution proposed in this paper. (1) “Signals Used” lists the physiological signals employed. Examples are Electrocardiogram (ECG), Heart Rate (HR), Galvanic Skin Response (GSR), Respiration (Resp.), and Electromyogram (EMG), among others. The choice of signals is crucial as it directly affects the accuracy and reliability of the detection system. Different signals provide varied information about an individual’s physiological state, impacting the performance of the employed models. (2) “Application” specifies whether the study focuses on stress (S) or emotion (E) detection, whereas the second considers nuanced mental conditions. (3) “Device Type” specifies the type of device used to collect the physiological signals, such as wearables, smartwatches, wristbands, GSR devices, or pulse oximetry. The device type influences the monitoring system’s accuracy, practicality, comfort, and user acceptance. Only some devices allow for seamless monitoring in real-world settings, which is especially valuable for frail people. (4) “Model Type” lists the machine learning or statistical models used for classification tasks. Notice how, except [3] and [48], all works rely on simple off-the-shelf machine learning models (SVM, kNN, and pLDA [^5] Extended Linear Discriminant Analysis.), which might not be suitable to detect emotions or stress levels that are not in the training data distribution, and therefore have limited capacity for generalization but also for personalization. In [3], an ensemble mechanism over Gradient Boosts (GB) is applied to classify stress levels by leveraging the expertise of multiple models. In this case, a linear regression meta-model is applied to the output of the GB base models to predict the outcome. In  [48], a clustering algorithm based on a Gaussian mixture convolutional variational autoencoder is employed to identify normal emotional patterns and detect anomalies. To the best of our knowledge, we are the first to use a state-of-the-art foundational model for time-series to tackle stress and abnormal emotion detection. (5) “Task” indicates whether the model’s task is classification (C) or anomaly detection (AD). Most works tackle stress and emotion monitoring as classification problems, categorizing states into predefined classes. The only exception is UAED [48], which performs an anomaly detection task. We argue that approaching stress monitoring as an anomaly detection problem is more principled and inherently more explainable than doing fine-grained emotion classification.
@@ 2.2 Anomaly Detection in Time Series
Following [48], we cast stress monitoring as an anomaly detection problem. We identified five methods proposed in the literature for anomaly detection in time series. Distance-based outlier detectors consider the distance of a point from its k-nearest neighbors. Density-based methods (among others [42, 38]) consider the density of the point and its neighbors. Prediction-based methods (among others, [6]) calculate the difference between the predicted and true values to detect anomalies. Reconstruction-based methods (among others [13]) compare the input signal and the reconstructed one in the output layer, typically using autoencoders. These methods assume anomalies are difficult to reconstruct and lost when the signal is mapped to lower dimensions, so a higher reconstruction error indicates a higher anomaly score.
GAN-based anomaly detection methods have gained popularity recently. MAD-GAN [28] combines the discriminator output with reconstruction error to detect anomalies in multivariate time series. TadGAN [15] uses a cycle-consistent GAN architecture with an encoder-decoder generator and proposes several ways to compute reconstruction error combined with critic outputs. HypAD [13] proposes using hyperbolic uncertainty for anomaly detection. TranAD [41] is a transformer-based model that uses adversarial training and self-conditioning to capture broader temporal patterns effectively. USAD [5] combines adversarial training with an autoencoder, amplifying reconstruction errors for anomalies while ensuring robust learning.
Gaussian Mixture Model (GMM) methods include DAGMM [49], which integrates a deep autoencoder with a GMM for joint optimization, enhancing the alignment of dimensionality reduction and density estimation for unsupervised anomaly detection. Zhu et al. [48] utilize a 2D Convolutional Variational Autoencoder with GMM latent distributions, transforming ECG signals into image-like formats for anomaly detection.
Graph-based methods have also emerged as viable approaches for time-series anomaly detection. MTAD-GAT [47] employs Graph Attention Networks (GAT) to model inter-feature and temporal dependencies, optimizing forecasting and reconstruction tasks jointly. GDN [12] learns sensor relationships via graph structure learning and uses attention mechanisms for detecting deviations from learned patterns.
Among other deep learning-based approaches, OmniAnomaly [38] employs stochastic latent variables to model normal patterns in multivariate time series. CAE-M [46] combines a convolutional autoencoder with a memory network for robust anomaly detection, incorporating a Maximum Mean Discrepancy (MMD) penalty to reduce noise impact. MSCRED [44] captures multi-scale system status using signature matrices and applies convolutional LSTM to model temporal dependencies, diagnosing anomalies based on inter-sensor correlations. LSTM-NDT [22] utilizes LSTMs to forecast telemetry values with a dynamic thresholding mechanism for anomaly detection.
@@ 3 Method
UniTS [14] is the first foundation model [^6] Foundation models are pre-trained on broad data at scale and can be adapted to a wide range of downstream tasks. for multivariate time series, designed to perform in a unified way various tasks such as forecasting, imputation, classification, and anomaly detection. It employs self-attention to capture relationships across both the sequence and variable dimensions. This is particularly beneficial in our multi-sensor domain since, if one variable has an anomaly, the model can assess how this anomaly correlates with the patterns observed in other variables, thereby improving detection accuracy. Furthermore, UniTS is designed to cohesively integrate signals of different types and lengths. To achieve this, it employs multi-scale processing techniques to handle time-series data of different lengths, which involves processing the data at various temporal resolutions.
@@ 3.1 Problem Definition and Preliminaries
For each time point $t$, an anomaly score $S_{t}=g(\mathbf{x}_{t},f(\mathbf{X}_{train}))$ is computed, where $g$ measures the deviation of $\mathbf{x}_{t}$ from the expected behavior defined by $f$. An observation $\mathbf{x}_{t}$ is classified as abnormal if its anomaly score $S_{t}$ exceeds a predefined threshold $\tau$. Thus, the set of detected anomalies is given by $\mathbf{A}=\{\mathbf{x}_{t}\mid S_{t}>\tau,t\in\{1,2,\ldots,T\}\}$.
@@
Figure 1: Fine-tuning of the UniTS architecture for anomaly detection. The multivariate time series in input gets transformed into several tokens, which are then passed through $N$ different blocks of the UniTS model. UniTS reconstructs the sample tokens, which are unpatched and compared against the real tokens. The comparison uses a dynamic threshold to discern between anomalies and normality. For visualization purposes, we do not illustrate the prompt tokens concatenated to the sample ones.
@@ 3.2 Framework Overview
UniTS – see Figure 1 – is a versatile multi-task model featuring a unified network architecture. Without loss of generality, we illustrate the fine-tuning process of UniTS for anomaly detection, specifically using multivariate time series data like heart rate (HR) and heart rate variability (HRV). The process begins with the input time series data, $\mathbf{X}$, which contains both HR and HRV signals. This data is split into smaller segments or ”tokens” passed to the UniTS model. The model consists of $N$ stacked blocks, each incorporating a modified transformer architecture. These blocks include key components like Time Multi-Head Self-Attention (Time MHSA), Variable MHSA, and Dynamic Feed-Forward Networks (Dynamic FFN). Gates control the flow of information between these components. Once processed by these blocks, the UniTS model outputs reconstructed tokens, represented as $\tilde{z}_{\mathbf{X}}$, which are then unfolded or unpatched into time series data. The reconstruction is compared to the original tokens using a dynamic threshold mechanism. This comparison helps identify anomalies in the time series. Normal HR and HRV signals are classified as “normality,” whereas significant deviations from the expected patterns are flagged as “anomalies,” as shown in the green and red segments. The dynamic threshold process ultimately discerns between normal and abnormal signals, visualizing anomalies for detection.
Input Tokenization
In UniTS, input tokenization is a foundational process that converts raw time series data into structured representations, enabling the model to process complex multivariate time series tasks effectively. Each token corresponds to either a temporal step or a variable, encapsulating temporal dynamics and inter-variable relationships. This tokenization scheme allows the model to apply attention mechanisms across time and variables, thereby improving its capacity to capture dependencies. The model avoids task-specific preprocessing by unifying inputs into tokens, ensuring generalization across forecasting, classification, imputation, and anomaly detection. The structured nature of tokenized inputs enhances adaptability, making it integral for multi-task learning. In more detail, UniTS introduces three [^7] Although we do not use the task token in our scenario, we illustrate how they can be integrated into the framework. We point the reader to [14] for more details. unique token types: i.e., sample, prompt, and task tokens, each fulfilling a specific role in time series analysis.
  • 1.
@@
  • 2.
    Prompt tokens. They are defined as learnable embeddings $z_{p}\in\mathbb{R}^{p\times n\times d}$ where $p$ is the number of tokens. These tokens incorporate the task UniTS needs to perform. Notice that, in Figure 1, we do not illustrate the prompt tokens proposed in the original paper since we only exploit the forecasting prompt in our anomaly detection scenario. Nevertheless, for completeness, we invite the reader to visualize that these prompt tokens are prepended to the input series tokens as in Eq. (2).
@@
  • 3.
    Task tokens. The UniTS framework comes with an additional type of tokens, namely the task ones. Since we want to reconstruct tokens to be able to detect anomalies, we do not explicitly rely on the task tokens as suggested in [14]. In this scenario, we want to compare the reconstructed tokens with those coming from the ground truth (GT) and pinpoint the anomalies along the series. In detail, instead of injecting task tokens into the input, we want to reconstruct the input tokens.
UniTS modules
The model takes in input $z_{\text{Anomaly}}$ and feeds it to $N$ blocks of modified transformer architecture to handle heterogeneous multi-domain data with varying dynamics and the number of variables. We use the original UniTS architecture, as shown in Figure 1. In more detail, we use Time and Variable Multi-Head Self-Attention (MHSA) blocks, a Dynamic Feed-Forward Network (FFN), and gating modules. We refer the reader to the original paper’s appendix [14] for more detail on the transformer and tower modules.
  • 1.
    Time and Variable MHSA. As suggested in the original paper, we use a two-way self-attention for feature and time dimensions. This approach contrasts with previous methods that apply self-attention to time or variable dimensions but not tensors. Time and variable self-attention effectively handle time series samples with various numbers of features $n$ and different time lengths $t$.
  • 2.
    Dynamic FFN. The transformer block is modified by incorporating a dynamic operator into an FFN layer. This modification enables the FFN to capture dependencies between tokens, unlike the standard FFN, which processes embedding vectors point-wise.
  • 3.